# Neutrino mixing parameters and masses from $\Delta(96)\rtimes H_{CP}$ in the tri-direct CP approach

### v1

@@ Appendix A Group theory of $\Delta(96)$
$\hskip-21.68121pt1C_{1}:\{1\}\,,$
$\hskip-21.68121pt3C_{2}:\{c^{2},d^{2},c^{2}d^{2}\},$
$\hskip-21.68121pt12C_{2}:\{ab,abc,abc^{2},abc^{3},a^{2}b,a^{2}bd,a^{2}bd^{2},a^{2}bd^{3},b,bcd,bc^{2}d^{2},bc^{3}d^{3}\},$
$\hskip-21.68121pt32C_{3}:\{a,ac,ac^{2},ac^{3},ad,ad^{2},ad^{3},acd,acd^{2},acd^{3},ac^{2}d,ac^{2}d^{2},ac^{2}d^{3},ac^{3}d,ac^{3}d^{2},ac^{3}d^{3},a^{2},a^{2}c,$
$\hskip-21.68121pt\qquad\quad\ a^{2}c^{2},a^{2}c^{3},a^{2}d,a^{2}d^{2},a^{2}d^{3},a^{2}cd,a^{2}cd^{2},a^{2}cd^{3},a^{2}c^{2}d,a^{2}c^{2}d^{2},a^{2}c^{2}d^{3},a^{2}c^{3}d,a^{2}c^{3}d^{2},a^{2}c^{3}d^{3}\},$
$\hskip-21.68121pt3C^{(1)}_{4}:\{cd^{2},cd^{3},c^{2}d^{3}\},$
$\hskip-21.68121pt3C^{(2)}_{4}:\{c^{2}d,c^{3}d,c^{3}d^{2}\},$
$\hskip-21.68121pt6C_{4}:\{c,d,cd,c^{3},d^{3},c^{3}d^{3}\},$
$\hskip-21.68121pt12C_{4}:\{abd^{2},abcd^{2},abc^{2}d^{2},abc^{3}d^{2},a^{2}bc^{2},a^{2}bc^{2}d,a^{2}bc^{2}d^{2},a^{2}bc^{2}d^{3},bc^{2},bc^{3}d,bd^{2},bcd^{3}\},$
$\hskip-21.68121pt12C^{(1)}_{8}:\{abd,abcd,abc^{2}d,abc^{3}d,a^{2}bc^{3},a^{2}bc^{3}d,a^{2}bc^{3}d^{2},a^{2}bc^{3}d^{3},bc,bc^{2}d,bc^{3}d^{2},bd^{3}\},$
$\hskip-21.68121pt12C^{(2)}_{8}:\{abd^{3},abcd^{3},abc^{2}d^{3},abc^{3}d^{3},a^{2}bc,a^{2}bcd,a^{2}bcd^{2},a^{2}bcd^{3},bc^{3},bd,bcd^{2},bc^{2}d^{3}\}\,,$ (A.2)
@@
[…]
@@
The number of irreducible representations of a group is equal to the number of its conjugacy class, therefore $\Delta(96)$ has ten irreducible representations: two one-dimensional representations $\bm{1_{m}}$, one two-dimensional representation $\bm{2}$, six three-dimensional representations $\bm{3_{m}}$ and $\bm{\bar{3}_{m}}$ and $\bm{\hat{3}_{m}}$ and one six-dimensional representation $\bm{6}$, where $m=0,1$. The representation matrices for the generators $a$, $b$, $c$ and $d$ in different irreducible representations are listed in table 6. It should be noted that the representations $\bm{\bar{3}_{m}}$ are complex conjugate representations of $\bm{3_{m}}$. Together with the representation $\bm{6}$, these triplet representations $\bm{3_{m}}$ and $\bm{\bar{3}_{m}}$ form faithful representations of $\Delta(96)$, while $\bm{\hat{3}_{m}}$ are not. Then the Kronecker products between different irreducible representations can be easily obtained:
$\hskip-21.68121pt\bm{1_{m}}\otimes\bm{1_{n}}=\bm{1_{[m+n]}},\quad\bm{1_{m}}\otimes\bm{2}=\bm{2},\quad\bm{1_{m}}\otimes\bm{3_{n}}=\bm{3_{[m+n]}},\quad\bm{1_{m}}\otimes\bm{\bar{3}_{n}}=\bm{\bar{3}_{[m+n]}},$
$\hskip-21.68121pt\bm{1_{m}}\otimes\bm{\hat{3}_{n}}=\bm{\hat{3}_{[m+n]}},\quad\bm{1_{m}}\otimes\bm{6}=\bm{6},\quad\bm{2}\otimes\bm{2}=\bm{1_{0}}\oplus\bm{1_{1}}\oplus\bm{2},\quad\bm{2}\otimes\bm{3_{m}}=\bm{3_{0}}\oplus\bm{3_{1}},$
$\hskip-21.68121pt\bm{2}\otimes\bm{\bar{3}_{m}}=\bm{\bar{3}_{0}}\oplus\bm{\bar{3}_{1}},\quad\bm{2}\otimes\bm{\hat{3}_{m}}=\bm{\hat{3}_{0}}\oplus\bm{\hat{3}_{1}},\quad\bm{2}\otimes\bm{6}=\bm{6_{i}}\oplus\bm{6_{ii}},\quad\bm{3_{m}}\otimes\bm{3_{n}}=\bm{\bar{3}_{0}}\oplus\bm{\bar{3}_{1}}\oplus\bm{\hat{3}_{[m+n+1]}},$
$\hskip-21.68121pt\bm{3_{m}}\otimes\bm{\bar{3}_{n}}=\bm{1_{[m+n]}}\oplus\bm{2}\oplus\bm{6},\quad\bm{3_{m}}\otimes\bm{\hat{3}_{n}}=\bm{\bar{3}_{[m+n+1]}}\oplus\bm{6},\quad\bm{3_{m}}\otimes\bm{6}=\bm{3_{0}}\oplus\bm{3_{1}}\oplus\bm{\hat{3}_{0}}\oplus\bm{\hat{3}_{1}}\oplus\bm{6},$
$\hskip-21.68121pt\bm{\bar{3}_{m}}\otimes\bm{\bar{3}_{n}}=\bm{3_{0}}\oplus\bm{3_{1}}\oplus\bm{\hat{3}_{[m+n+1]}},~{}~{}\bm{\bar{3}_{m}}\otimes\bm{\hat{3}_{n}}=\bm{3_{[m+n+1]}}\oplus\bm{6},~{}~{}\bm{\bar{3}_{m}}\otimes\bm{6}=\bm{\bar{3}_{0}}\oplus\bm{\bar{3}_{1}}\oplus\bm{\hat{3}_{0}}\oplus\bm{\hat{3}_{1}}\oplus\bm{6},$
$\hskip-21.68121pt\bm{\hat{3}_{m}}\otimes\bm{\hat{3}_{n}}=\bm{1_{[m+n]}}\oplus\bm{2}\oplus\bm{\hat{3}_{0}}\oplus\bm{\hat{3}_{1}},\quad\bm{\hat{3}_{m}}\otimes\bm{6}=\bm{3_{0}}\oplus\bm{3_{1}}\oplus\bm{\bar{3}_{0}}\oplus\bm{\bar{3}_{1}}\oplus\bm{6},$
$\hskip-21.68121pt\bm{6}\otimes\bm{6}=\bm{1_{0}}\oplus\bm{1_{1}}\oplus\bm{2_{S}}\oplus\bm{2_{A}}\oplus\bm{3_{0}}\oplus\bm{3_{1}}\oplus\bm{\bar{3}_{0}}\oplus\bm{\bar{3}_{1}}\oplus\bm{\hat{3}_{0}}\oplus\bm{\hat{3}_{1}}\oplus\bm{6_{S}}\oplus\bm{6_{A}}\,,$ (A.8)
@@
where $m,n=0,1$, and we have defined $[m+n]\equiv m+n$ (mod 2). The symbol $\bm{6_{i}}$ and $\bm{6_{ii}}$ stand for two six-dimensional representations $\bm{6}$ that appear in the Kronecker products. The subscript “$S$” (“$A$”) refers to symmetric (antisymmetric) combinations. In the following, we list the CG coefficients in our basis. The notation $\alpha_{i}$ ($\beta_{i}$) are denoted the elements of the first (second) representation. Furthermore, we shall adopt the following notations facilitate the expressions of CG coefficients
@@
$\bm{1_{m}}$ $1$ $(-1)^{m}$ $1$ $1$
$\bm{2}$ $-\frac{1}{2}\begin{pmatrix}1&\sqrt{3}\\ -\sqrt{3}&1\\ \end{pmatrix}$ $\begin{pmatrix}1&0\\ 0&-1\\ \end{pmatrix}$ $\mathbb{1}_{2}$ $\mathbb{1}_{2}$
@@
$\begin{array}[]{l}\bm{\hat{3}_{[m+n]}}:~{}\begin{pmatrix}\alpha_{1}\beta_{1}\\ \alpha_{1}\beta_{2}\\ \alpha_{1}\beta_{3}\end{pmatrix}\end{array}$ $\begin{array}[]{l}\bm{6}:~{}P_{6}^{m}\begin{pmatrix}\alpha_{1}\beta_{1}\\ \alpha_{1}\beta_{2}\\ \alpha_{1}\beta_{3}\\ \alpha_{1}\beta_{4}\\ \alpha_{1}\beta_{5}\\ \alpha_{1}\beta_{6}\end{pmatrix}\\ \end{array}$ $\begin{array}[]{l}\bm{1_{0}}:\alpha_{1}\beta_{1}+\alpha_{2}\beta_{2}\\ \bm{1_{1}}:\alpha_{1}\beta_{2}-\alpha_{2}\beta_{1}\\ \bm{2}:\begin{pmatrix}\alpha_{1}\beta_{1}-\alpha_{2}\beta_{2}\\ -\alpha_{1}\beta_{2}-\alpha_{2}\beta_{1}\\ \end{pmatrix}\end{array}$ $\begin{array}[]{l}\bm{3_{m}}:\quad~{}~{}\begin{pmatrix}(\alpha_{1}-\sqrt{3}\alpha_{2})\beta_{1}\\ -2\alpha_{1}\beta_{2}\\ (\alpha_{1}+\sqrt{3}\alpha_{2})\beta_{3}\end{pmatrix}\\ \bm{3_{[m+1]}}:\begin{pmatrix}(\alpha_{2}+\sqrt{3}\alpha_{1})\beta_{1}\\ -2\alpha_{2}\beta_{2}\\ (\alpha_{2}-\sqrt{3}\alpha_{1})\beta_{3}\end{pmatrix}\end{array}$
@@
$\begin{array}[]{l}\bm{\bar{3}_{m}}:\quad~{}~{}\begin{pmatrix}(\alpha_{1}-\sqrt{3}\alpha_{2})\beta_{1}\\ -2\alpha_{1}\beta_{2}\\ (\alpha_{1}+\sqrt{3}\alpha_{2})\beta_{3}\end{pmatrix}\\ \bm{\bar{3}_{[m+1]}}:\begin{pmatrix}(\alpha_{2}+\sqrt{3}\alpha_{1})\beta_{1}\\ -2\alpha_{2}\beta_{2}\\ (\alpha_{2}-\sqrt{3}\alpha_{1})\beta_{3}\end{pmatrix}\end{array}$ $\begin{array}[]{l}\bm{\hat{3}_{m}}:\quad~{}~{}\begin{pmatrix}(\alpha_{1}-\sqrt{3}\alpha_{2})\beta_{1}\\ -2\alpha_{1}\beta_{2}\\ (\alpha_{1}+\sqrt{3}\alpha_{2})\beta_{3}\end{pmatrix}\\ \bm{\hat{3}_{[m+1]}}:\begin{pmatrix}(\alpha_{2}+\sqrt{3}\alpha_{1})\beta_{1}\\ -2\alpha_{2}\beta_{2}\\ (\alpha_{2}-\sqrt{3}\alpha_{1})\beta_{3}\end{pmatrix}\end{array}$ $\begin{array}[]{l}\bm{\bar{3}_{[m+n]}}:\quad~{}\begin{pmatrix}\alpha_{2}\beta_{3}-\alpha_{3}\beta_{2}\\ \alpha_{3}\beta_{1}-\alpha_{1}\beta_{3}\\ \alpha_{1}\beta_{2}-\alpha_{2}\beta_{1}\end{pmatrix}\\ \bm{\bar{3}_{[m+n+1]}}:\begin{pmatrix}\alpha_{2}\beta_{3}+\alpha_{3}\beta_{2}\\ \alpha_{3}\beta_{1}+\alpha_{1}\beta_{3}\\ \alpha_{1}\beta_{2}+\alpha_{2}\beta_{1}\end{pmatrix}\\ \bm{\hat{3}_{[m+n+1]}}:\begin{pmatrix}\alpha_{1}\beta_{1}\\ \alpha_{2}\beta_{2}\\ \alpha_{3}\beta_{3}\end{pmatrix}\end{array}$
@@
$\begin{array}[]{l}\bm{3_{[n+m]}}:\quad~{}\begin{pmatrix}\alpha_{2}\beta_{3}-\alpha_{3}\beta_{2}\\ \alpha_{3}\beta_{1}-\alpha_{1}\beta_{3}\\ \alpha_{1}\beta_{2}-\alpha_{2}\beta_{1}\end{pmatrix}\\ \bm{3_{[n+m+1]}}:\begin{pmatrix}\alpha_{2}\beta_{3}+\alpha_{3}\beta_{2}\\ \alpha_{3}\beta_{1}+\alpha_{1}\beta_{3}\\ \alpha_{1}\beta_{2}+\alpha_{2}\beta_{1}\end{pmatrix}\\ \bm{\hat{3}_{[m+n+1]}}:\begin{pmatrix}\alpha_{1}\beta_{1}\\ \alpha_{2}\beta_{2}\\ \alpha_{3}\beta_{3}\end{pmatrix}\end{array}$ $\begin{array}[]{l}\bm{1_{[m+n]}}:~{}~{}\alpha_{1}\beta_{1}+\alpha_{2}\beta_{2}+\alpha_{3}\beta_{3}\\ \bm{2}:P_{2}^{m+n}\begin{pmatrix}\alpha_{1}\beta_{1}-2\alpha_{2}\beta_{2}+\alpha_{3}\beta_{3}\\ \sqrt{3}(\alpha_{3}\beta_{3}-\alpha_{1}\beta_{1})\\ \end{pmatrix}\\ \bm{\hat{\bm{6}}}:P_{6}^{m+n}\begin{pmatrix}\alpha_{1}\beta_{3}\\ \alpha_{2}\beta_{1}\\ \alpha_{3}\beta_{2}\\ \alpha_{3}\beta_{1}\\ \alpha_{2}\beta_{3}\\ \alpha_{1}\beta_{2}\\ \end{pmatrix}\end{array}$ $\begin{array}[]{l}\bm{\bar{3}_{[m+n+1]}}:\begin{pmatrix}\alpha_{1}\beta_{1}\\ \alpha_{2}\beta_{2}\\ \alpha_{3}\beta_{3}\end{pmatrix}\\ \bm{6}:P_{6}^{m+n}\begin{pmatrix}\alpha_{2}\beta_{1}\\ \alpha_{3}\beta_{2}\\ \alpha_{1}\beta_{3}\\ \alpha_{2}\beta_{3}\\ \alpha_{1}\beta_{2}\\ \alpha_{3}\beta_{1}\end{pmatrix}\end{array}$

### v2

@@ Appendix A Group theory of $\Delta(96)$
$\thinspace 1C_{1}:\{1\}\,$,
$\thinspace 3C_{2}:\{c^{2},d^{2},c^{2}d^{2}\},$
$\thinspace 12C_{2}:\{ab,abc,abc^{2},abc^{3},a^{2}b,a^{2}bd,a^{2}bd^{2},a^{2}bd^{3},b,bcd,bc^{2}d^{2},bc^{3}d^{3}\},$
$\thinspace 32C_{3}:\{a,ac,ac^{2},ac^{3},ad,ad^{2},ad^{3},acd,acd^{2},acd^{3},ac^{2}d,ac^{2}d^{2},ac^{2}d^{3},ac^{3}d,ac^{3}d^{2},ac^{3}d^{3},a^{2},a^{2}c,$
$\thinspace\qquad\quad\ a^{2}c^{2},a^{2}c^{3},a^{2}d,a^{2}d^{2},a^{2}d^{3},a^{2}cd,a^{2}cd^{2},a^{2}cd^{3},a^{2}c^{2}d,a^{2}c^{2}d^{2},a^{2}c^{2}d^{3},a^{2}c^{3}d,a^{2}c^{3}d^{2},a^{2}c^{3}d^{3}\},$
$\thinspace 3C^{(1)}_{4}:\{cd^{2},cd^{3},c^{2}d^{3}\},$
$\thinspace 3C^{(2)}_{4}:\{c^{2}d,c^{3}d,c^{3}d^{2}\},$
$\thinspace 6C_{4}:\{c,d,cd,c^{3},d^{3},c^{3}d^{3}\},$
$\thinspace 12C_{4}:\{abd^{2},abcd^{2},abc^{2}d^{2},abc^{3}d^{2},a^{2}bc^{2},a^{2}bc^{2}d,a^{2}bc^{2}d^{2},a^{2}bc^{2}d^{3},bc^{2},bc^{3}d,bd^{2},bcd^{3}\},$
$\thinspace 12C^{(1)}_{8}:\{abd,abcd,abc^{2}d,abc^{3}d,a^{2}bc^{3},a^{2}bc^{3}d,a^{2}bc^{3}d^{2},a^{2}bc^{3}d^{3},bc,bc^{2}d,bc^{3}d^{2},bd^{3}\},$
$\thinspace 12C^{(2)}_{8}:\{abd^{3},abcd^{3},abc^{2}d^{3},abc^{3}d^{3},a^{2}bc,a^{2}bcd,a^{2}bcd^{2},a^{2}bcd^{3},bc^{3},bd,bcd^{2},bc^{2}d^{3}\}\,$, (A.2)
@@
    $\begin{aligned} &Z_{2}^{c^{2}}=\left\{1,c^{2}\right\},&&Z_{2}^{d^{2}}=\left\{1,d^{2}\right\},&&Z_{2}^{c^{2}d^{2}}=\left\{1,c^{2}d^{2}\right\},\\ &Z_{2}^{ab}=\left\{1,ab\right\},&&Z_{2}^{abc}=\left\{1,abc\right\},&&Z_{2}^{abc^{2}}=\left\{1,abc^{2}\right\},&&Z_{2}^{abc^{3}}=\left\{1,abc^{3}\right\},\\ &Z_{2}^{a^{2}b}=\left\{1,a^{2}b\right\},&&Z_{2}^{a^{2}bd}=\left\{1,a^{2}bd\right\},&&Z_{2}^{a^{2}bd^{2}}=\left\{1,a^{2}bd^{2}\right\},&&Z_{2}^{a^{2}bd^{3}}=\left\{1,a^{2}bd^{3}\right\},\\ &Z_{2}^{b}=\left\{1,b\right\},&&Z_{2}^{bcd}=\left\{1,bcd\right\},&&Z_{2}^{bc^{2}d^{2}}=\left\{1,bc^{2}d^{2}\right\},&&Z_{2}^{bc^{3}d^{3}}=\left\{1,bc^{3}d^{3}\right\}\,.\end{aligned}$ (A.3)
@@
The number of irreducible representations of a group is equal to the number of its conjugacy class. Therefore $\Delta(96)$ has ten irreducible representations: two one-dimensional representations $\bm{1_{m}}$, one two-dimensional representation $\bm{2}$, six three-dimensional representations $\bm{3_{m}}$ and $\bm{\bar{3}_{m}}$ and $\bm{\hat{3}_{m}}$ and one six-dimensional representation $\bm{6}$, where $m=0,1$. The representation matrices for the generators $a$, $b$, $c$ and $d$ in different irreducible representations are listed in table 6. It should be noted that the representations $\bm{\bar{3}_{m}}$ are complex conjugate representations of $\bm{3_{m}}$. Together with the representation $\bm{6}$, these triplet representations $\bm{3_{m}}$ and $\bm{\bar{3}_{m}}$ form faithful representations of $\Delta(96)$, while $\bm{\hat{3}_{m}}$ are not. Then the Kronecker products between different irreducible representations can be easily obtained:
$\thinspace\bm{1_{m}}\otimes\bm{1_{n}}=\bm{1_{[m+n]}},\quad\bm{1_{m}}\otimes\bm{2}=\bm{2},\quad\bm{1_{m}}\otimes\bm{3_{n}}=\bm{3_{[m+n]}},\quad\bm{1_{m}}\otimes\bm{\bar{3}_{n}}=\bm{\bar{3}_{[m+n]}},$
$\thinspace\bm{1_{m}}\otimes\bm{\hat{3}_{n}}=\bm{\hat{3}_{[m+n]}},\quad\bm{1_{m}}\otimes\bm{6}=\bm{6},\quad\bm{2}\otimes\bm{2}=\bm{1_{0}}\oplus\bm{1_{1}}\oplus\bm{2},\quad\bm{2}\otimes\bm{3_{m}}=\bm{3_{0}}\oplus\bm{3_{1}},$
$\thinspace\bm{2}\otimes\bm{\bar{3}_{m}}=\bm{\bar{3}_{0}}\oplus\bm{\bar{3}_{1}},\quad\bm{2}\otimes\bm{\hat{3}_{m}}=\bm{\hat{3}_{0}}\oplus\bm{\hat{3}_{1}},\quad\bm{2}\otimes\bm{6}=\bm{6_{i}}\oplus\bm{6_{ii}},\quad\bm{3_{m}}\otimes\bm{3_{n}}=\bm{\bar{3}_{0}}\oplus\bm{\bar{3}_{1}}\oplus\bm{\hat{3}_{[m+n+1]}},$
$\thinspace\bm{3_{m}}\otimes\bm{\bar{3}_{n}}=\bm{1_{[m+n]}}\oplus\bm{2}\oplus\bm{6},\quad\bm{3_{m}}\otimes\bm{\hat{3}_{n}}=\bm{\bar{3}_{[m+n+1]}}\oplus\bm{6},\quad\bm{3_{m}}\otimes\bm{6}=\bm{3_{0}}\oplus\bm{3_{1}}\oplus\bm{\hat{3}_{0}}\oplus\bm{\hat{3}_{1}}\oplus\bm{6},$
$\thinspace\bm{\bar{3}_{m}}\otimes\bm{\bar{3}_{n}}=\bm{3_{0}}\oplus\bm{3_{1}}\oplus\bm{\hat{3}_{[m+n+1]}},~~\bm{\bar{3}_{m}}\otimes\bm{\hat{3}_{n}}=\bm{3_{[m+n+1]}}\oplus\bm{6},~~\bm{\bar{3}_{m}}\otimes\bm{6}=\bm{\bar{3}_{0}}\oplus\bm{\bar{3}_{1}}\oplus\bm{\hat{3}_{0}}\oplus\bm{\hat{3}_{1}}\oplus\bm{6},$
$\thinspace\bm{\hat{3}_{m}}\otimes\bm{\hat{3}_{n}}=\bm{1_{[m+n]}}\oplus\bm{2}\oplus\bm{\hat{3}_{0}}\oplus\bm{\hat{3}_{1}},\quad\bm{\hat{3}_{m}}\otimes\bm{6}=\bm{3_{0}}\oplus\bm{3_{1}}\oplus\bm{\bar{3}_{0}}\oplus\bm{\bar{3}_{1}}\oplus\bm{6},$
$\thinspace\bm{6}\otimes\bm{6}=\bm{1_{0}}\oplus\bm{1_{1}}\oplus\bm{2_{S}}\oplus\bm{2_{A}}\oplus\bm{3_{0}}\oplus\bm{3_{1}}\oplus\bm{\bar{3}_{0}}\oplus\bm{\bar{3}_{1}}\oplus\bm{\hat{3}_{0}}\oplus\bm{\hat{3}_{1}}\oplus\bm{6_{S}}\oplus\bm{6_{A}}\,$, (A.8)
@@
where $m,n=0,1$, and we have defined $m+n]\equiv m+n$ (mod 2). The symbol $\bm{6_{i}}$ and $\bm{6_{ii}}$ stand for two six-dimensional representations $\bm{6}$ that appear in the Kronecker products. The subscript “SS” (“AA”) refers to symmetric (antisymmetric) combinations. In the following, we list the CG coefficients in our basis. The notation $\alpha_{i}$ ($\beta_{i}$) are denoted the elements of the first (second) representation. Furthermore, we shall adopt the following notations facilitating the expressions of CG coefficients
@@
$\bm{1_{m}}$ 11 $(-1)^{m}$ 11 11
$\bm{2}$ $\frac{1}{2}\begin{pmatrix}1&\sqrt{3}\\ -\sqrt{3}&1\\ \end{pmatrix}$ $\begin{pmatrix}1&0\\ 0&-1\\ \end{pmatrix}$ $\mathbb{1}_{2}$ $\mathbb{1}_{2}$
@@
$\begin{array}[]{l}\bm{\hat{3}_{[m+n]}}:~\begin{pmatrix}\alpha_{1}\beta_{1}\\ \alpha_{1}\beta_{2}\\ \alpha_{1}\beta_{3}\end{pmatrix}\end{array}$ $\begin{array}[]{l}\bm{6}:~P_{6}^{m}\begin{pmatrix}\alpha_{1}\beta_{1}\\ \alpha_{1}\beta_{2}\\ \alpha_{1}\beta_{3}\\ \alpha_{1}\beta_{4}\\ \alpha_{1}\beta_{5}\\ \alpha_{1}\beta_{6}\end{pmatrix}\\ \end{array}$ $\begin{array}[]{l}\bm{1_{0}}:\alpha_{1}\beta_{1}+\alpha_{2}\beta_{2}\\ \bm{1_{1}}:\alpha_{1}\beta_{2}-\alpha_{2}\beta_{1}\\ \bm{2}:\begin{pmatrix}\alpha_{1}\beta_{1}-\alpha_{2}\beta_{2}\\ -\alpha_{1}\beta_{2}-\alpha_{2}\beta_{1}\\ \end{pmatrix}\end{array}$ $\begin{array}[]{l}\bm{3_{m}}:\qquad~~\begin{pmatrix}(\alpha_{1}-\sqrt{3}\alpha_{2})\beta_{1}\\ -2\alpha_{1}\beta_{2}\\ (\alpha_{1}+\sqrt{3}\alpha_{2})\beta_{3}\end{pmatrix}\\ \bm{3_{[m+1]}}:\begin{pmatrix}(\alpha_{2}+\sqrt{3}\alpha_{1})\beta_{1}\\ -2\alpha_{2}\beta_{2}\\ (\alpha_{2}-\sqrt{3}\alpha_{1})\beta_{3}\end{pmatrix}\end{array}$
@@
$\begin{array}[]{l}\bm{\bar{3}_{m}}:\qquad~~\begin{pmatrix}(\alpha_{1}-\sqrt{3}\alpha_{2})\beta_{1}\\ -2\alpha_{1}\beta_{2}\\ (\alpha_{1}+\sqrt{3}\alpha_{2})\beta_{3}\end{pmatrix}\\ \bm{\bar{3}_{[m+1]}}:\begin{pmatrix}(\alpha_{2}+\sqrt{3}\alpha_{1})\beta_{1}\\ -2\alpha_{2}\beta_{2}\\ (\alpha_{2}-\sqrt{3}\alpha_{1})\beta_{3}\end{pmatrix}\end{array}$ $\begin{array}[]{l}\bm{\hat{3}_{m}}:\qquad~~\begin{pmatrix}(\alpha_{1}-\sqrt{3}\alpha_{2})\beta_{1}\\ -2\alpha_{1}\beta_{2}\\ (\alpha_{1}+\sqrt{3}\alpha_{2})\beta_{3}\end{pmatrix}\\ \bm{\hat{3}_{[m+1]}}:\begin{pmatrix}(\alpha_{2}+\sqrt{3}\alpha_{1})\beta_{1}\\ -2\alpha_{2}\beta_{2}\\ (\alpha_{2}-\sqrt{3}\alpha_{1})\beta_{3}\end{pmatrix}\end{array}$ $\begin{array}[]{l}\bm{\bar{3}_{[m+n]}}:\qquad~\begin{pmatrix}\alpha_{2}\beta_{3}-\alpha_{3}\beta_{2}\\ \alpha_{3}\beta_{1}-\alpha_{1}\beta_{3}\\ \alpha_{1}\beta_{2}-\alpha_{2}\beta_{1}\end{pmatrix}\\ \bm{\bar{3}_{[m+n+1]}}:\begin{pmatrix}\alpha_{2}\beta_{3}+\alpha_{3}\beta_{2}\\ \alpha_{3}\beta_{1}+\alpha_{1}\beta_{3}\\ \alpha_{1}\beta_{2}+\alpha_{2}\beta_{1}\end{pmatrix}\\ \bm{\hat{3}_{[m+n+1]}}:\begin{pmatrix}\alpha_{1}\beta_{1}\\ \alpha_{2}\beta_{2}\\ \alpha_{3}\beta_{3}\end{pmatrix}\end{array}$
@@
$\begin{array}[]{l}\bm{3_{[n+m]}}:\qquad~\begin{pmatrix}\alpha_{2}\beta_{3}-\alpha_{3}\beta_{2}\\ \alpha_{3}\beta_{1}-\alpha_{1}\beta_{3}\\ \alpha_{1}\beta_{2}-\alpha_{2}\beta_{1}\end{pmatrix}\\ \bm{3_{[n+m+1]}}:\begin{pmatrix}\alpha_{2}\beta_{3}+\alpha_{3}\beta_{2}\\ \alpha_{3}\beta_{1}+\alpha_{1}\beta_{3}\\ \alpha_{1}\beta_{2}+\alpha_{2}\beta_{1}\end{pmatrix}\\ \bm{\hat{3}_{[m+n+1]}}:\begin{pmatrix}\alpha_{1}\beta_{1}\\ \alpha_{2}\beta_{2}\\ \alpha_{3}\beta_{3}\end{pmatrix}\end{array}$ $\begin{array}[]{l}\bm{1_{[m+n]}}:~~\alpha_{1}\beta_{1}+\alpha_{2}\beta_{2}+\alpha_{3}\beta_{3}\\ \bm{2}:P_{2}^{m+n}\begin{pmatrix}\alpha_{1}\beta_{1}-2\alpha_{2}\beta_{2}+\alpha_{3}\beta_{3}\\ \sqrt{3}(\alpha_{3}\beta_{3}-\alpha_{1}\beta_{1})\\ \end{pmatrix}\\ \bm{\hat{\bm{6}}}:P_{6}^{m+n}\begin{pmatrix}\alpha_{1}\beta_{3}\\ \alpha_{2}\beta_{1}\\ \alpha_{3}\beta_{2}\\ \alpha_{3}\beta_{1}\\ \alpha_{2}\beta_{3}\\ \alpha_{1}\beta_{2}\\ \end{pmatrix}\end{array}$ $\begin{array}[]{l}\bm{\bar{3}_{[m+n+1]}}:\begin{pmatrix}\alpha_{1}\beta_{1}\\ \alpha_{2}\beta_{2}\\ \alpha_{3}\beta_{3}\end{pmatrix}\\ \bm{6}:P_{6}^{m+n}\begin{pmatrix}\alpha_{2}\beta_{1}\\ \alpha_{3}\beta_{2}\\ \alpha_{1}\beta_{3}\\ \alpha_{2}\beta_{3}\\ \alpha_{1}\beta_{2}\\ \alpha_{3}\beta_{1}\end{pmatrix}\end{array}$